OpenStudy (annjewell1998):

Could you please help me solve this? I'm not looking for just the answer i'd like to know how to solve it please :) Mary had a meal the Carrington Cafe, which cost $19.70. If the tax was 5.5 %, how much did she pay in tax? Round to the nearest cent. a$1.08 b$5.50 c$1.97 d$2.20

OpenStudy (jango_in_dtown):

tax=(5.5/100*19.70)=1.0835 when rounded gives 1.08

OpenStudy (jango_in_dtown):

x% of P=(x/100*P)

OpenStudy (annjewell1998):

Thank you so much! This helped! :)